OpenStudy (calculusxy):

Medal !! A large container is being filled with water at the rate of 2 3/4 gallons per minute. If the container leaks at the rate of 1/8 gallons per minute, how long would it take to fill a 14-gallon container?

OpenStudy (calculusxy):

@Mertsj

OpenStudy (mertsj):

Let x = number of minutes to fill the container \[2.75x-.125x=14\]

OpenStudy (wolf1728):

It is filling at 2 3/4 - 1/8 = 2 5/8 gallons per minute. It is a 14 gallon container so we are solving for Time. Here's a handy calcualtor: http://www.1728.org/flow.htm So time = 5.3333 minutes

OpenStudy (calculusxy):

How do you calculate time?

OpenStudy (wolf1728):

That calculator shoud really be used to check your work, so I'll work that out the long way.

OpenStudy (wolf1728):

2.625 gallons per minute container is 14 gallons time = 14 gallons / 2.625 gallons / minute time = 14/2.625 minutes = 5.3333333333 minutes

OpenStudy (calculusxy):

Thank you. Can I also divide fractions to also compute time?

OpenStudy (wolf1728):

Just see what I posted above your posting.

OpenStudy (calculusxy):

Ok. But thanks.

OpenStudy (wolf1728):

But you do see it right? (The calculator is really handy to chekc the work though) :-)